A ball rolls down a hill with constant acceleration. After 2.0 seconds the ball has travelled 1.2 m, how far will it have travelled after 4.0 seconds?
1
Expert's answer
2020-01-07T12:46:15-0500

As per the given question,


Intial velocity is not given, so i am considering it zero,

hence applying the second law of motion,

"s= ut+\dfrac{at^2}{2}"

"1.2=0+\dfrac{a\times 2\times 2}{2}"

"a=\dfrac{1.2}{2}=0.6 m/sec^2"

Now, again applying the second law of motion,

"s_1=ut+\dfrac{at_2^2}{2}=0+\dfrac{0.6\times 4\times 4}{2}=4.8 m"

After the next 4 sec ball will travel

"S_3-S_1=\dfrac{a(t_2^2-t_1^2)}{2}=\dfrac{0.6\times (36-4)}{2}=9.6 m"